Computing and Interpreting Percent Depreciated and PPE Turnover The following footnote is from Note 8 to the 2018 of Tesla, IQuestion text Computing and Interpreting Percent Depreciated and PPE Turnover The following footnote is from Note 8 to the 2018 of Tesla, Inc.: Note 8- Property, Plant and Equipment Our property, plant and equipment, net, consisted of the following (in thousands): 2018 2017 Machinery, equipment, vehicles and office furniture $ 6,328,966 $ 4,251,711 Tooling 1,397,514 1,255,952 Leasehold improvements 960,971 789,751 Land and buildings 4,047,006 2,517,247 Computer equipment, hardware and software 487,421 395,067 Construction in progress 807,297 2,541,588 14,029,175 11,751,316 Less: Accumulated depreciation (2,699,098) (1,723,794) Total $ 11,330,077 $ 10,027,522 The summary of significant accounting policies included the following description of Tesla’s depreciation policies: Property, plant and equipment, including leasehold improvements, are recognized at cost less accumulated depreciation. Depreciation is generally computed using the straight-line method over the estimated useful lives of the respective assets, as follows: Machinery, equipment, vehicles and office furniture 2 to 12 years Building and building improvements 15 to 30 years Computer equipment and software 3 to 10 years Depreciation for tooling is computed using the units-of-production method whereby capitalized costs are amortized over the total estimated productive life of the respective assets. As of December 31, 2018, the estimated productive life for Model S and Model X tooling was 325,000 vehicles based on our current estimates of production. As of December 31, 2018, the estimated productive life for Model 3 tooling was 1,000,000 vehicles based on our current estimates of production. 1. Tesla’s revenue totaled $21,461,268 ($ thousands) in 2018. Compute its PPE turnover for the year. Do not round until your final answer. Round final answer to one decimal place. Answer times 2. Compute the percent depreciated ratio for 2018. Round answer to one decimal place (ex: 0.2345 = 23.5%). Answer %
Answer #1

Part 1

PPE turnover = revenue / average PPE = 21461268/((11330077+10027522)/2) = 2.01

Part 2

Percent depreciated ratio for 2018 =Accumulated depreciation / Depreciable asset cost = 2699098/(14029175-807297) = 20.4%